Resolve the concentration of 10 mL of 0.1 N hydrochloric acid accidentally added with 90 mL of distilled water. Answer with solution please. Thank you.


1
Expert's answer
2021-10-10T10:25:29-0400

Solution:


C1 = 0.1 N

V1 = 10 mL

C2 = unknown

V2 = V1 + V(H2O) = 10 mL + 90 mL = 100 mL


Hence,

(0.1 N) × (10 mL) = C2 × (100 mL)

C2 = (0.1 N) × (10 mL) / (100 mL) = 0.01 N

C2 = 0.01 N


Answer: The concentration of hydrochloric acid will be 0.01 N